# Securing WiFi Fingerprint-based Indoor Localization Systems from Malicious Access Points

### v1

@@ V-B2 Comparative Analysis of AAIFU and Proposed Schemes
False positive rate (FPR): Figure 5(a) shows that the FPR ranges between 2% 6% and 2% 4% for AAIFU and proposed schemes, respectively. Due to the temporal effect both schemes exhibit FPR.
@@ V-C Impact of Malicious Effect Mitigation
Figure. 6 shows that the curve for 50% malicious AP exhibits significant degradation from the curve with no malicious AP across all types of attacks. The proposed scheme, which includes database update and model retraining, shows an almost unnoticeable deviation from the curve with no malicious AP for every type of attack. However, the proposed scheme that does not include the database updation and model retraining exhibits degraded performance compared to the proposed scheme with the database updation curve. By updating the database and retraining the model after identifying malicious APs, the localization error is minimized. The AAIFU scheme [15] focuses on the steady change in the behavior of APs rather than the dynamic malicious effects during the online phase. This scheme produces results similar to our proposed scheme without requiring the database update, as shown in Fig. 6. The CDF curve for the AAIFU scheme is smoother for the use of a regressor and the CDF curve for the proposed scheme appears more like a staircase for the use of a classifier.
It is observed from experiments that the proposed scheme effectively handles at most 60% malicious APs. After that the performance deviates to a greater extent.
@@ V-D Execution Time
The proposed scheme exhibits a gradual increase in runtime with the increasing number of malicious APs as shown in Table V. The AAIFU scheme randomly detects malicious APs, which is displayed by the non-gradual increase of runtime. The AAIFU scheme spends a significant amount of time in the detection phase than the proposed scheme as the proposed scheme uses only some simple statistical calculations whereas the AAIFU scheme uses GBDTRegressors. The AAIFU scheme detects fewer APs than the proposed scheme. Hence, even though both schemes use regressors for updating database, the proposed scheme takes more time than the AAIFU scheme. The proposed scheme takes more time than the AAIFU scheme in model retraining, indicating that training a classifier is more time-consuming than a regressor. Lastly, in the inference phase, the proposed scheme involves regressors to predict the values of malicious APs in online queries, resulting in greater time consumption compared to the AAIFU scheme. The overall time consumed by the proposed scheme is approximately 94% less than the AAIFU scheme.
@@ VI Conclusion
We have proposed a reliable and efficient long-term indoor localization system by incorporating malicious AP detection and effect mitigation modules. Experimental results demonstrate that the proposed scheme achieves detection accuracy above 95% for all attack types. Besides, the malicious effect mitigation module ensures the accurate functioning of the indoor localization system by restoring the system’s performance to the initial state where the system operates without any malicious APs. In addition, the proposed indoor localization system cuts down the execution time by approximately 94% compared to existing works. We believe the proposed scheme enhances the existing literature by offering a faster and more reliable indoor localization system. We plan to extend this work by considering device diversity and attack scenarios for online and offline databases.

### v2

@@ V-B2 Comparative Analysis of AAIFU and Proposed Schemes
False positive rate (FPR): Figure 5(a) shows that the FPR ranges between 2%6% and 2%4% for AAIFU and proposed schemes, respectively. Due to the temporal effect, both schemes exhibit FPR.
@@ V-C Impact of Malicious Effect Mitigation
Figure. 6 shows that the curve for 50% malicious AP exhibits significant degradation from the curve with no malicious AP across all types of attacks. The proposed scheme, which includes database update and model retraining, shows an almost unnoticeable deviation from the curve with no malicious AP for every type of attack. However, the proposed scheme that does not include the database updation and model retraining exhibits degraded performance compared to the proposed scheme with the database updation curve. By updating the database and retraining the model after identifying malicious APs, the localization error is minimized. The AAIFU scheme [15] focuses on the steady change in the behavior of APs rather than the dynamic malicious effects during the online phase. This scheme produces results similar to our proposed scheme without requiring the database update, as shown in Fig. 6. The CDF curve for the AAIFU scheme is smoother for the use of a regressor, and the CDF curve for the proposed scheme appears more like a staircase for the use of a classifier.
It is observed from experiments that the proposed scheme effectively handles at most 60% malicious APs. After that, the performance deviates to a greater extent.
@@ V-D Execution Time
The proposed scheme exhibits a gradual increase in runtime with the increasing number of malicious APs, as shown in Table V. The AAIFU scheme randomly detects malicious APs, which is displayed by the non-gradual increase of runtime. The AAIFU scheme spends a significant amount of time in the detection phase than the proposed scheme, as the proposed scheme uses only some simple statistical calculations, whereas the AAIFU scheme uses GBDTRegressors. The AAIFU scheme detects fewer APs than the proposed scheme. Hence, even though both schemes use regressors for updating database, the proposed scheme takes more time than the AAIFU scheme. The proposed scheme takes more time than the AAIFU scheme in model retraining, indicating that training a classifier is more time-consuming than a regressor. Lastly, in the inference phase, the proposed scheme involves regressors to predict the values of malicious APs in online queries, resulting in greater time consumption compared to the AAIFU scheme. The overall time consumed by the proposed scheme is approximately 94% less than the AAIFU scheme.
@@ VI Conclusion
We have proposed a reliable and efficient long-term indoor localization system by incorporating malicious AP detection and effect mitigation modules. Experimental results show that the proposed scheme achieves detection accuracy above 95% for all attack types. The malicious effect mitigation module ensures the accurate functioning of the indoor localization system by restoring the system’s performance to the initial state, where the system operates without any malicious APs. Besides, the proposed indoor localization system cuts down the execution time by approximately 94% compared to existing works. We believe the proposed scheme enhances the existing literature by offering a faster and more reliable indoor localization system. We plan to extend this work by considering device diversity, attack scenarios for online and offline databases, system scalability, and generalizability.